OpenStudy (anonymous):

The graph of f passes through (-9, 6) and is perpendicular to the line that has a x-int of 7 and y-int of -14 Find slope intercept

OpenStudy (anonymous):

does this just mean (-9, 6) (7, -14)

OpenStudy (anonymous):

Nope that point and the intercepts are two complete different lines You need to use the intercepts to develop a line equation and then using the slope of it figure out the slope of the equation you need and develop that one. I will help you one step at a time

OpenStudy (anonymous):

So the intercepts are giving you a set of two points (7,0) and (0,-14) we need to find a slope from these two points so \[\frac{ y2-y1 }{ x2-x1 } = \frac{ 0-(-14) }{ 0-7 }\] This equals -2 right?

OpenStudy (anonymous):

woops would be positive 2 since 14/7 to keep things consistent

OpenStudy (anonymous):

\[\frac{ 0-(-14) }{ 7-0 }\]

OpenStudy (anonymous):

With me so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok so this is our m so far for thisl ine we know that y=2x + b Well we don't need the b of this line since we know the slope A slope perpendicular to this line would be the opposite reciprocal so -1/2 so we have y= -1/2x + b and now we use (-9,6) to solve for b 6 = -1/2(9) + b 6= -9/2 + b 12/2 = -9/2 + b 17/2 = bSo our line perpendicular to the one formed with the intercepts is y = -1/2x + 17/2

OpenStudy (anonymous):

its all the knowing when to flip stuff that always confuses me. thank you again

OpenStudy (anonymous):

although webwork says the answer is wrong

OpenStudy (anonymous):

hmmm try y = 1/2x + 3/2

OpenStudy (anonymous):

yep :) right

OpenStudy (anonymous):

-1/2 though

OpenStudy (anonymous):

Ohhhh haha i copied the point wrong as 9 not -9

OpenStudy (anonymous):

6 = -1/2(-9) + b 6 = 9/2 + b 12/2 = 9/2 + b b = 3/2

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

Yep! Starting to get it more?